the magnetic susceptibility of a matter is 25*10^-6 . calculate its absolute and relation permeabilities
 

Grade:12th pass

1 Answers

Arun
5 years ago
Relative Permeability mu(r)= 1 + K/mu(0)
 
 = 1 + 25*10^-6/4*pi*10^-7 = 20.96
 
Absolute permeability = 
mu = mu(0) * mu(r)